A man is 3 times as old as his son. In 10 years he will be twice as old as his son will be. How old are they now?
Answer by stanbon(75887)

A man is 3 times as old as his son. In 10 years he will be twice as old as his son will be. How old are they now?
-----------------
Equations:
f = 3s
f+10 = 2(s+10)
-----
Substitute for "f" and solve for "s"::
3s+10 = 2s+20
s = 10 yrs (son's age now)
----
f = 3s = 30 yrs (father's age now)
-----------
Cheers
Stan H.
------------